# Region-aware Grasp Framework with Normalized Grasp Space for 6-DoF Grasping in Cluttered Scene

### v1

@@ IV-B3 Characteristics of Normalized Grasp Space
  • SE(2)-rotation invariance for ($\beta,\gamma$) and equivariance for $\theta$: When the patch $\boldsymbol{P}_{i}$ rotates along the z-axis in the camera frame with $\Delta\theta$, the grasps inside the patch rotates equivariantly:
    $\boldsymbol{G}^{*}_{t}=f(\boldsymbol{P}^{*}_{t})$ $=f(\mathbf{R}_{z}(\Delta\theta)\boldsymbol{P}^{*})$ (8)
    $\mathbf{R}^{*}_{t}=Rot(\boldsymbol{G}^{*}_{t})$ $=\mathbf{R}_{z}(\Delta\theta)Rot(\boldsymbol{G}_{t})$
    $=\mathbf{R}_{z}(\Delta\theta)\mathbf{R}^{*}$
    $=\mathbf{R}_{z}(\Delta\theta)\mathbf{R}_{z}(\theta)\mathbf{R}_{x}(\beta)\mathbf{R}_{y}(\gamma)$
    $=\mathbf{R}_{z}(\theta+\Delta\theta)\mathbf{R}_{x}(\beta)\mathbf{R}_{y}(\gamma).$ (9)
    SE(2)-rotation invariance and equivariance of Normalized Grasping Space reduce the difficulty of grasp model training and enhance robustness to different object poses and camera viewpoints.
@@
  • Scale-invariance: With the coordinates (XYZ) of the patches or the scene scaled by a specific factor $a\in\mathbb{R}^{+}$, benefiting from the normalization process, NGS produces the grasps with unchanged width and offset if the gripper width for normalization is scaled by the same factor $a$:
    $\boldsymbol{G}^{*}_{t}=f(\boldsymbol{P}^{*}_{t})$ $=f(Norm(\text{scale}(\boldsymbol{P},a)))$
    $=f(Norm(a\boldsymbol{P}|w_{gripper}=aw_{0}))$
    $=f(\frac{a(\boldsymbol{P}-\mathbf{p}_{i})}{aw_{0}})$
    $=f(Norm(\boldsymbol{P}|w_{gripper}=w_{0}))$
    $=f(\boldsymbol{P}^{*})=\boldsymbol{G}^{*}.$ (10)
    By adapting gripper width for different scenes or conducting multi-scale gripper inference, our proposed Normalized Grasping Space and grasp detection method can be directly applied to novel scenes with objects in different scales without any extra model training.
@@ IV-B4 Patch scale randomization
Above we consider the ideal scenario for applying Normalized Grasp Space. However, when detecting grasp in the real world, noise from depth sensors is non-negligible and may cause unstable patch extraction results with inaccurate size. Furthermore, considering the model generalization to grippers with different scales, the gripper sizes in grasp detection datasets are usually fixed. This means that though we train our model through a normalized scheme, it actually senses the regions with a relatively fixed receptive field size. Thus, during training, we add randomization to the patch receptive field $r_{i}$, that is, the final patch receptive field is randomly selected in the range $[\frac{r_{i}}{2},2*r_{i}]$, which can improve the robustness and generalization capability of our grasp network.
IV-C Regional Normalized Grasp Network
Based on Normalized Grasp Space, Regional Normalized Grasp Network processes normalized patches aiming at predicting normalized grasp attributes within these regional patches. Utilizing a coordinate-gated CNN, Region-aware Patch Feature Extraction extracts patch-wise grasp-relevant features fed into the latter Regional Multi-Grasp Generator to generate robust and consistent grasps in each local patch.
Figure 5: Detailed structure of Regional Normalized Grasp Network (RNGNet).
IV-C1 Region-aware Patch Feature Extraction
@@
To detect high-quality grasps in regions, our approach initially extracts patch-wise features in a region-aware manner. Given the grasp generation challenge within the Normalized Grasp Space, the 3D XYZ coordinates of the patches are crucial for understanding object contact and collision. Hence, it is imperative to incorporate not only the depth channel but also the overall 3D XYZ coordinates for grasp-relevant and region-aware feature extraction. For geometric feature extraction from 3D XYZ coordinates, most pointcloud-based methods utilize point-wise shared MLP combined with point sampling and grouping [32, 33, 25], which is effective but introduces a significant computation overhead. Thus, combining the advantages of pointcloud-based and CNN-based methods, as depicted in Fig. 5, we design an efficient network based on point-wise shared MLP and 2D convolution to process the normalized patches and extract patch features.
For coordinate-relevant tasks, such as grasp detection in the local patches, every pixel in the patch has different importance based on the distance to the patch center and its 3D position, which is highly relevant to object contact or collision. However, traditional CNN treats each pixel in the same way, ignoring the influence of the coordinates. As is discussed in former works [20, 15], regular 2D convolution networks are not adept at coordinate-relevant tasks. Similar to the tasks in [2], due to the effective normalization in NGS, grasp detection in the local patches is also layout-specific to a certain extent. A regular convolution kernel shares the same filter weights across the whole image, while the distance between nearby pixels may vary due to different depths to the camera.
@@
Thus, inspired by the gated CNN proposed in [13], as shown in the right bottom in Fig. 5, we consider integrating XYZ coordinate sensing by point-wise shared MLP. We establish a multi-stage spatially-varying convolution framework featuring a lightweight unit called PosGate. This unit, operating on 3D point coordinate XYZ maps, increases the dimensions of XYZ maps in different resolutions to obtain spatial feature gate values in each convolution stage. By multiplying the original feature maps with the gate values, PosGate enables selective amplification or attenuation of filters within convolution layers. By leveraging 6-channel RGBXYZ maps as input combined with multi-stage PosGates, RNGNet adeptly extracts patch features in a region-aware fashion.
IV-C2 Regional Multi-Grasp Generator
Similar to typical regional anchor-based grasp generator in [9, 40, 3], with the region-aware features extracted, we predefine a series of non-uniform rotation anchors generated via the Anchor Shifting algorithm in [3]. However, different from them, our Regional Multi-Grasp Generator focuses on detect grasps only using the patch features without any scene-level priors. Then, we consider the grasp detection as a multi-label classification on the three Euler angles while the normalized grasp translations and widths are regressed for each possible rotation anchor. All classification and regression heads are implemented utilizing separate MLPs (Multi-layer Perceptron).
Though similar in grasp generation pipeline, different from the two-part grasp representation introduced in [3] and the global-to-local grasp detection scheme [9, 40], our detected grasp poses are generated only depending on the region-aware feature of the patches. Thus, we ensure that consistent detection results will be generated for identical input patches. In addition, such a region-aware framework shows better generalizability to novel scenes and robustness to input noises.
IV-D Loss
$L=L_{\theta\_cls}+a*L_{\theta\_reg}+b*L_{\gamma\beta}+c*L_{t}+d*L_{w},$ (11)
IV-E Implemetation Details
In the patch extraction, we extract normalized patches with size $S\times S=64\times 64$ by default. We process the input image as the same size $640\times 360$ and adopt the pre-trained Grasp Heatmap Model checkpoint in [3] for a fair performance comparison. In the feature extraction part, we directly adopt the ResNet-18 [12] with half network width as the mainstream of our backbone in RNGNet. For the implementation of the PosGate module, a 3-layer shared MLP (Multi-layer Perceptron) with the same width as the corresponding feature maps is adopted to generate gate value for feature maps of each resolution. In our multi-grasp generator, the overall settings about rotation anchor are kept unchanged from those in the [3]. AdamW optimizer is adopted to train our RNGNet. More hyperparameters are listed in the Supplementary Materials.
Figure 6: Pipeline and visualization of Normalized Patch Dataset Preparation.
TABLE I: Results on GraspNet Dataset, showing APs on RealSense/Kinect split and method time usage
@@ IV-E Implemetation Details
Method Seen $\uparrow$ Similar $\uparrow$ Novel $\uparrow$ Average $\uparrow$ Paras $\downarrow$ Time1/ms $\downarrow$
GPD [39] 22.87 / 24.38 21.33 / 23.18 8.24 / 9.58 17.48 / 19.05 - -
PointnetGPD [18] 25.96 / 27.59 22.68 / 24.38  9.23 / 10.66 19.29 / 20.88 - -
GraspNet-baseline [9] 27.56 / 29.88 26.11 / 27.84 10.55 / 11.51 21.41 / 23.08 - 126
TransGrasp [22] 39.81 / 35.97 29.32 / 29.71 13.83 / 11.41 27.65 / 25.70 - -
HGGD [3] 58.35 / 56.85 47.93 / 43.93 22.10 / 18.59 42.79 / 39.79 3.42M 34
Scale Balanced Grasp [24] 58.95 /   - 52.97 /   - 22.63 /   - 44.85 /   - - 242
GSNet [40] 65.70 / 61.19 53.75 / 47.39 24.31 / 19.01 47.92 / 42.53 15.4M 61
GRE-Grasp [4]   -  / 65.19   -  / 54.09   -  / 22.29   -  / 47.19 23.3M -
AnyGrasp w/ CD [10] 66.12 /   - 56.09 /   - 24.81 /   - 49.01 /   - 24.7M 198
RNGNet 75.20 / 72.23 66.62 / 58.43 32.38 / 26.05 58.06 / 52.24 3.66M 17
RNGNet w/ CD 76.28 / 72.89 68.26 / 59.42 32.84 / 26.12 59.13 / 52.81 3.66M 20
@@
“-”: Result Unavailable, CD:Collision Detection, the grasp post-processing algorithm proposed in [9] and utilized by [10].
1 Including data processing, network inference, and grasp post-processing. All results are evaluated with $batch\_size=1$ on Ubuntu20.04 with AMD 5600x CPU and a single NVIDIA RTX 3060Ti GPU.
Tested with the provided codebases.
Reimplemented with official codebase.
More detailed results are shown in the Supplementary Materials.
V Dataset Experiments
V-A Normalized Patch Dataset Preparation
Currently, most grasp datasets focus on providing object-level grasp labels [8, 27] or scene-level grasp label [9, 45], which are not directly suitable for our normalized scheme training. Considering the downstream application on the real robot platform and data capacity among all the datasets, GraspNet-1Billion [9] is a large-scale grasp dataset widely adopted in recent 6-DoF grasp detection research [10, 3, 24, 4], which provides large-scale training data and a standard evaluation platform for the task of general robotic grasping. The total dataset contains 97,280 real-world RGBD images with over one billion grasp poses and in total 190 cluttered scenes, which are split into 100 training scenes and 90 testing scenes (30 with seen objects, 30 with similar objects, and 30 with novel objects).
Although GraspNet-1Billion offers a large number of RGBD images and grasps annotations, based on the proposed Normalized Grasp Space, we need to construct a patch-based grasp dataset providing normalized patches and grasps annotations. As is shown in Fig. 6, utilizing the scene-level annotations provided in the GraspNet-1Billion dataset, we generate a new patch-based grasp dataset for Regional Normalized Grasp Network training. Firstly we obtain the object segmentation mask in the original dataset and adopt a Gaussian kernel to dilate the mask image for covering possible graspable regions as much as possible. Then a fixed number of patch center candidates are randomly sampled on the dilated mask image and cropped. Also, as is mentioned in [34], depth information is important to generate collision-free grasp detection results. Thus, we add domain randomization to the patch center depth on the RGBD image planes, which provides a patch-based dataset covering more of the possible circumstances and improves the robustness of our training method.
V-B Evaluation Metrics
With the large-scale patch-based grasp dataset, we can fully train our RNGNet for grasp detection inside the local patches. Then we evaluate our method on the 90 testing scenes (30 with seen objects, 30 with similar objects and 30 with novel objects) of the GraspNet-1Billion dataset. During our experiments, we follow the official evaluation pipeline and code of the GraspNet-1Billion dataset, in which detected grasp poses are first filtered with non-maximum suppression and then the top 50 grasp poses are evaluated with force-closure [26] metrics under a series of friction coefficients condition. In our experiments, we evaluate our method on all three parts of the test set and adopt the AP (Average Precision), which indicates the overall average precision along all given friction coefficients and better displays the general performance of each method.
V-C Quantitative Results
In our quantitative experiments, we compare the overall performance of our RNGNet with multiple typical grasp detection algorithms trained on the same dataset training split, including GPD [39], PointnetGPD [18], GraspNet-Baseline [9], GSNet [40], TransGrasp [22], Scale Balanced Grasp [24], HGGD [3] and Grasp Region Exploration [4]. As Table. I indicates, RNGNet achieves excellent performance, reaching an average 58.06/52.24 AP by RealSense/Kinect cameras. Compared with other methods, RNGNet outperforms current state-of-the-art by a large margin on all dataset splits, especially on the unseen (similar and novel) splits, proving the efficacy and genelizabilty of our proposed region-aware grasp framework. Furthermore, considering the computation cost for one shot of grasp detection, our method also achieves the fastest speed with the best grasp detection quality, which implies the potential of our approach to be applied in dynamic scenes or on widely used computation-resource-constrained platforms in robot deployment.
As shown in Fig. 7, utilizing the force-closure metric, we conduct grasp evaluation under different difficulties with different object friction coefficients. With lower object friction, the object is more likely to drop from the gripper and thus requires more high-quality grasp detection results. It can be clearly seen in the figure that RNGNet surpasses former SOTAs by a large margin with different object friction coefficients, especially for the settings with lower friction coefficients. These results further prove the robustness and generalizability of our proposed region-aware framework.
Figure 7: Grasp detection performance (mAP) on the GraspNet-1B dataset comparison under different difficulties (friction coefficient).
@@ V-C Quantitative Results
Figure 8: Qualitative results on GraspNet-1Billion dataset compared to HGGD and GSNet, covering three parts of the test set (seen/similar/novel). Top 50 grasps after grasp-NMS [9] are displayed. Color implies the network-predicted grasp confidence (red: high, blue: low).
V-D Qualitative Results
@@
In addition to the quantitative results, we also conduct grasp detection visualization with HGGD [3] and GSNet [40] on three parts of the test set (seen: scene100-129, similar: scene130-159, novel: scene160-189). As depicted in Fig. 8, RNGNet exhibits superior grasp detection quality and achieves a higher grasp coverage rate in cluttered scenes than prior works, even in scenarios involving occlusion or partial observation. This enhancement provides the robot with more feasible action alternatives, benefiting scene-level grasp planning and execution. Furthermore, when compared with [3, 10], it is noteworthy that our proposed region-aware framework enables the generation of grasps better aligned with object centers and surfaces, consequently enhancing grasp stability. The overall qualitative results underscore the validity of our representation and approach in cluttered scenes, offering a more comprehensive perspective.
TABLE II: Ablation Experiments of Normalized Grasp Space, Showing APs on Realsense Split
@@
V-E Ablation Studies
V-E1 Normalized Grasp Space
Firstly, we perform ablation studies on our proposed Normalized Grasp Space, which primarily consists of three components: Patch Normalization, Depth-adaptive Patch Extraction, and Scale Randomization. The results in Table. II illustrate the effects of these three components on Normalized Grasp Space. As expected, both Patch Normalization and Depth-adaptives Patch Extraction significantly enhance performance across all scenes, proving the effectiveness of the novel viewpoint about generating grasps in the region-aware and grasp-centric spaces. Scale Randomization also improves the performance of our methods on the dataset.
TABLE III: Ablation Experiments of Input Modalities and Network Structure, Showing APs on Realsense Split
@@ V-E1 Normalized Grasp Space
TABLE IV: Ablation Experiments of Backbone Selection, Showing APs on Realsense Split
Backbone Size Seen Similar Novel Time1
SparseCNN [5] $64\times 64$ 72.17 61.90 29.85 74
$32\times 32$ 71.76 61.63 29.89 29
$16\times 16$ 68.16 56.68 25.76 18
RNGNet $64\times 64$ 75.20 66.62 32.38 8
$32\times 32$ 72.13 62.81 29.86 5
$16\times 16$ 67.04 55.35 25.63 4
1 Tested with the same setting mentioned above. Results in milliseconds (ms), including only network inference time.
TABLE V: Ablation Experiments of Patch Localization Algorithm, Showing APs on Realsense Split
Localization Seen Similar Novel Time1 (ms)
Heatmap 75.20 66.62 32.38 17
Foreground 74.75 66.06 33.59 30
1 Same as in Table. 2.
TABLE VI: Ablation Experiments of Region-aware Grasp Framework, Showing APs on Realsense Split
Distance Boundary Seen Similar Novel
0.1 70.02 59.89 29.88
0.2 75.20 66.62 32.38
0.4 69.60 62.77 28.08
0.8 35.50 31.16 13.10
V-E2 Region Normalized Grasp Network
We conduct ablation experiments from three perspectives to evaluate the efficiency and robustness of our RNGNet.
Network Structure: Table. III presents the ablation experiments conducted on the RNGNet with different input modalities. Remarkably, even with only Z maps as input, it already achieves a commendable performance. The introduction of XY maps or RGB images as supplementary inputs leads to considerable improvements. While geometric information from Z maps plays a crucial role in 6-DoF grasp detection, the additional positional information from XY maps and the inclusion of colors contribute to enhanced feature extraction and are especially beneficial for generalization to previously unseen (similar and novel) scenes. Finally, the application of PosGate to feature maps in all stages yields further improvements across all test splits, with a negligible increase in the number of parameters (approximately 36k, or about 1% of the total parameters). This supports our assumption regarding the influence of camera perspective models and layout-specific features in Normalized Grasp Space to some extent.
Backbone Selection: To further prove the efficiency and robustness of our 2D-convolution-based network, in Table. IV, we replace the backbone with the widely used [40, 4, 23] high-dimensional sparse convolution backbone, MinkowskiFCNN [5]. As is shown in the results, our simple 2D convolution network is more efficient than the more complicated sparse convolution network in experiments, especially with high-resolution input. When facing low-resolution input, SparseCNN performs slightly better due to its particular design for sparse input. Significantly, both RNGNet and SparseCNN show quite consistent performance with input patches of different resolutions, which proves the robustness of our region-aware framework.
Patch Localization: We also consider the influence of the patch localization method. In Table. V, we change the introduced heatmap-based patch center selection method to simply sampling more centers on filtered object point cloud by removing the table point cloud (foreground sampling). On the one hand, heatmap assists RNGNet to concentrate more on the critical part in the scene and get a faster inference speed. On the other hand, RNGNet is robust to inaccurate patch centers input with a simple model-free foreground sampling. The higher performance of foreground sampling method on the novel split of the test set further proves the generalization capability of our region-aware framework.
V-E3 Region-aware Grasp Setting
Additionally, we conduct ablation studies on our proposed region-aware grasp setting which constrain grasp detection close to the patch centers by adjusting the regional grasp center distance boundary in Equation (6). A large distance boundary means that we allow the method to detect grasps far away from the grasp patch center, while a small one means we apply stricter constraints on the region-aware scheme. To prove the effectiveness of our Region-aware Grasp Framework, we generate new datasets and train extra models under different levels of restrictions on the grasp center point. As shown in Table VI, restricting the grasp to be located close to the grasp with a reasonable distance shows optimal results on all the dataset splits, enhancing robustness to clutter and occlusions and improving generalization across environments. In our region-aware framework, each patch extracted focuses on generating feasible grasps within, and regional grasps from multiple patches compose the scene-level grasp prediction results.
TABLE VII: Results of Table-top Clutter Scene Clearance Experiments
Scene Objects RNGNet HGGD[3] AnyGrasp[10]
@@ V-E3 Region-aware Grasp Setting
Figure 9: Real-world settings: (a) System setup for our physical grasping experiments. (b) Test object samples in the grasping trails.
VI Realworld Experiments
We first conduct real-world experiments on clutter scene clearance under more challenging conditions, comparing the outcomes with those of two competitors. In addition to the core framework tailored for grasp detection in cluttered scenes, we leverage the adaptability of the approach to extend our framework for multi-scale grasp detection by simply rescaling $w_{gripper}$ and enabling closed-loop grasping in dynamic scenes by real-time grasp generation.
VI-A Cluttered Scene Clearance

### v2

@@ 3.1 Normalized Grasp Space
2) SE(2)-rotation invariance for ($\beta,\gamma$) and equivariance for $\theta$: When the patch $\boldsymbol{P}_{i}$ rotates along the z-axis in the camera frame with $\Delta\theta$, the grasps inside the patch rotates equivariantly:
$\mathbf{R}^{*}_{\mathcal{T}}=\mathbf{R}_{z}(\Delta\theta)\mathbf{R}^{*}=\mathbf{R}_{z}(\Delta\theta)\mathbf{R}_{z}(\theta)\mathbf{R}_{x}(\beta)\mathbf{R}_{y}(\gamma)=\mathbf{R}_{z}(\theta+\Delta\theta)\mathbf{R}_{x}(\beta)\mathbf{R}_{y}(\gamma).$ (5)
@@
3) Scale-invariance: With the coordinates (XYZ) of the patches or the scene scaled by a factor $a\in\mathbb{R}^{+}$, we get the same normalized patch $\boldsymbol{P}^{*}_{i}$ and the grasps inside the patch keep invariant if the receptive field for normalization is scaled by the same factor $a$:
$\boldsymbol{G}^{*}_{\mathcal{T}}=f(\boldsymbol{P}^{*}_{\mathcal{T}})=f(\text{Norm}(a\boldsymbol{P}|w_{ref}=aw_{0}))=f(\frac{a(\boldsymbol{P}-\mathbf{p}_{i})}{aw_{0}})=f(\boldsymbol{P}^{*})=\boldsymbol{G}^{*}.$ (6)
@@
Patch Scale Randomization: Above, we consider the ideal scenario for applying Normalized Grasp Space. However, in the real world, noise from depth sensors is non-negligible and may cause unstable patch extraction. Furthermore, the gripper sizes in grasp detection datasets are usually fixed, which means that though we train our model through a normalized scheme, it senses the regions with a fixed receptive field. Thus, we add randomization to the receptive field $w_{ref}$ during training. The receptive field is randomly selected in the range $[\frac{w_{ref}}{2},2*w_{ref}]$, which can improve the robustness and generalization capability of our grasp network.
3.2 Regional Normalized Grasp Network
Figure 4: Detailed structure of RNGNet.
@@ 3.2 Regional Normalized Grasp Network
Region-aware Patch Feature Extraction: To efficiently detect grasps in regions, our approach extracts patch-wise features in a region-aware manner. The XYZ coordinates of the patches are crucial for understanding object contact and collision. Hence, it is imperative to incorporate the overall 3D XYZ coordinates instead of only the depth channel for feature extraction. For geometric feature extraction from point-wise coordinates, previous methods utilize shared MLP (Multi-layer Perceptron) combined with sampling and grouping [22, 12, 23], which is effective but introduces a significant computation overhead. Thus, combining the advantages of point-wise shared MLPs and CNNs, as depicted in Fig. 4, we design an efficient network to process the normalized patches.
@@
Inspired by the gated CNN in [24, 25], we establish a multi-stage Gated CNN with a novel Point-wise Gate (PG) Module. Point-wise Gate Module is built with shared MLP operating on 3D point coordinate XYZ maps, increasing the dimensions of XYZ maps in different resolutions to obtain spatial feature gate values in each convolution stage. By multiplying the original feature maps with the gate values, Point-wise Gate enables selective amplification or attenuation of filters within convolution layers.
Rotation Heatmap Predictor: Based on the regional anchor-based grasp generator in [3, 6, 15], with the region-aware features extracted, we predefine a series of non-uniform rotation anchors for $(\gamma,\beta)$ generated via the Anchor Shifting algorithm in [15], whose Cartesian product forms a 2D rotation heatmap (one axis for $\gamma$, the other of $\beta$). Then, the grasp detection can be considered a two-class semantic segmentation problem, while the normalized grasp widths and extra center offsets are regressed for each possible rotation (pixel in the heatmap).
@@
Method Average $\uparrow$ Seen $\uparrow$ Similar $\uparrow$ Novel $\uparrow$ Paras $\downarrow$ Time1/ms $\downarrow$
GPD [10] 17.48 / 19.05 22.87 / 24.38 21.33 / 23.18 8.24 / 9.58 - -
PointnetGPD [11] 19.29 / 20.88 25.96 / 27.59 22.68 / 24.38  9.23 / 10.66 - -
GraspNet-baseline [3] 21.41 / 23.08 27.56 / 29.88 26.11 / 27.84 10.55 / 11.51 - 126
TransGrasp [26] 27.65 / 25.70 39.81 / 35.97 29.32 / 29.71 13.83 / 11.41 - -
HGGD [15] 42.79 / 39.79 58.35 / 56.85 47.93 / 43.93 22.10 / 18.59 3.42M 34
Scale Balanced Grasp [8] 44.85 /   - 58.95 /   - 52.97 /   - 22.63 /   - - 242
GSNet [6] 47.92 / 42.53 65.70 / 61.19 53.75 / 47.39 24.31 / 19.01 15.4M 61
AnyGrasp w/ CD [9] 49.01 /   - 66.12 /   - 56.09 /   - 24.81 /   - 24.7M 198
RNGNet 58.06 / 52.24 75.20 / 72.23 66.62 / 58.43 32.38 / 26.05 3.66M 17
RNGNet w/ CD 59.13 / 52.81 76.28 / 72.89 68.26 / 59.42 32.84 / 26.12 3.66M 20
@@
“-”: Result Unavailable; CD:Collision Detection, the grasp post-processing algorithm proposed in [3] and utilized by [9].
1 Evaluated with $batch\_size=1$ on Ubuntu20.04 with AMD 5600x CPU and a single NVIDIA RTX 3060Ti GPU.
Reimplemented or tested with the provided codebases.   More detailed results are shown in Table 7 of the Appendices.
Table 1: Results on GraspNet dataset. Showing APs on RealSense/Kinect split and model efficiency.
@@
Figure 5: Qualitative results covering seen/similar/novel test set. Top 50 grasps after grasp-NMS [3] are displayed. Color implies the predicted grasp confidence (red: high, blue: low).
Figure 6: Performance (Average AP) curve. Tested under different difficulties (friction coefficient).
4 Dataset Experiments
Dataset and Metrics: GraspNet-1Billion [3] is a large-scale grasp dataset widely adopted in recent 6-DoF grasp detection research [9, 15, 8, 19], which provides large-scale training data and a standard evaluation platform for the task of general robotic grasping. During our experiments, we follow the official evaluation pipeline and code of the GraspNet-1Billion dataset, in which detected grasp poses are first filtered with non-maximum suppression, and then the top 50 grasp poses are evaluated with force-closure [27] metrics under a series of friction coefficients condition.
Quantitative Results: We compare the overall performance of our RNGNet with multiple typical grasp detection algorithms trained on the same dataset training split. As Table 1 indicates, RNGNet achieves excellent performance, reaching an average 58.06/52.24 AP by RealSense/Kinect cameras. Compared with other methods, RNGNet outperforms current state-of-the-art by a large margin on all dataset splits, especially on the unseen (similar and novel) splits, proving the efficacy and generalizability of our proposed region-aware grasp framework. Furthermore, considering the computation cost for one shot of grasp detection, our method achieves the fastest speed with the best grasp detection quality, which implies the potential of our approach to be applied in dynamic scenes or on widely used computation-resource-constrained platforms in robot deployment.
As shown in Fig. 6, we conduct grasp evaluation using the force-closure metric under different difficulties with different object friction coefficients. With lower object friction, the object is more likely to drop from the gripper and thus requires more high-quality grasp detection results. It can be clearly seen in the figure that RNGNet surpasses former SOTAs by a large margin with different object friction coefficients, especially for the settings with lower friction coefficients.
As is indicated in Table 2, following the setting in [8], we evaluate the grasp quality at all scales (S: Small, M: Medium, L: Large). RNGNet significantly improves performance with all the metrics, especially in the small-scale grasps and the novel split, proving the proposed NGS’s effectiveness. RNGNet can learn more robust and generalizable region features across different scales with NGS.
Method Seen Similar Novel
APS APM APL APS APM APL APS APM APL
Scale Balanced Grasp [8] 13.47 48.12 61.81 6.23 37.90 53.89 7.60 17.04 23.10
Scale Balanced Grasp + OBS [8] 18.29 52.60 64.34 10.03 42.77 57.09 9.29 18.74 24.36
RNGNet 22.94 64.06 68.50 19.34 59.55 64.58 21.51 29.84 28.57
Table 2: Multi-scale results. Grasps with different scales are tested separately on RealSense split. OBS: Object Balanced Sampling proposed in [8] which requires extra instance segmentation.
@@ 4 Dataset Experiments
Qualitative Results: We also conduct grasp detection visualization with HGGD [15] and GSNet [6] on three parts of the test set (seen: scene100-129, similar: scene130-159, novel: scene160-189). As depicted in Fig. 5, RNGNet exhibits superior grasp detection quality and achieves a higher grasp coverage rate in cluttered scenes than prior works, even in scenarios involving occlusion or partial observation. This enhancement provides the robot with more feasible action alternatives, benefiting scene-level grasp planning and execution. Furthermore, when compared with [15, 9], it is noteworthy that our proposed region-aware framework enables the generation of grasps better aligned with object centers and surfaces, consequently enhancing grasp stability. The overall qualitative results underscore the validity of our representation and approach in cluttered scenes, offering a more comprehensive perspective.
Normalized Grasp Space Ablation: We perform ablation studies on our proposed Normalized Grasp Space, which primarily consists of Patch Normalization, Depth-adaptive Patch Extraction, and Scale Randomization. The results in Table 3 illustrate the effects of these three components on Normalized Grasp Space. As expected, both Patch Normalization and Depth-adaptive Patch Extraction significantly enhance performance across all scenes, proving the effectiveness of the novel viewpoint about generating grasps in the region-aware and grasp-centric spaces. Additionally, Scale Randomization also boosts the performance of our methods on the dataset.
Region Normalized Grasp Network Ablation: Table 4 presents the ablation experiments conducted on the RNGNet with different input modalities. Remarkably, our method achieves a commendable performance even with only depth maps as input. Introducing XY maps or RGB images as supplementary inputs leads to considerable improvements. While geometric information from Z maps plays a crucial role in 6-DoF grasp detection, the additional positional information from XY maps and the inclusion of colors contribute to enhanced feature extraction and are especially beneficial for generalization to previously unseen (similar and novel) scenes. Finally, applying Point-wise Gate Module to feature maps in all stages yields further improvements across all test splits, with a negligible increase in the number of parameters (approximately 36k, or about 1% of the total parameters). These findings support the efficacy of our approach, which is built with a combination of point-based networks and CNNs.
@@
Table 3: Normalized Grasp Space ablation. Tested on Realsense Split.
@@
PG: Point-wise Gate Module
Table 4: Input modality and network ablation. Tested on Realsense Split.
Figure 7: Robot settings and object samples.
Scene Objects RNGNet HGGD AnyGrasp
@@
Table 5: Real-world clutter clearance results. : failed to grasp all the objects.
5 Realworld Experiments